# Conformal Depression Prediction

### v1

@@ III-B Depression Recognition with UQ
The straightforward method for uncertainty quantification in depression recognition is to utilize a DNN to estimate the conditional probability distribution $\mathcal{P}(y|X=x)$ of depressed individuals. Given our lack of knowledge regarding the true $\mathcal{P}(y|X=x)$, currently, the most common approach assumes that depression target $y$ follow a Gaussian distribution $y\sim\mathcal{G}(\hat{y},\sigma^{2})$. Using the negative log-likelihood (NLL) for the Gaussian distribution [46], we can derive the loss function:
$\mathcal{L}_{NLL}=\sum_{i=1}^{N}\frac{log\sigma_{i}^{2}}{2}+\frac{(y_{i}-\hat{y}_{i})^{2}}{2\sigma_{i}^{2}}$ (2)
@@
Minimizing Eq 2 as the learning objective during training allows us to obtain a DNN capable of predicting the distribution $\mathcal{P}(y|X=x)$ for any given individual. Furthermore, in the context of the Gaussian distribution, the quantile can be calculated as follows:
$\hat{y}_{\alpha}=\hat{y}+\sigma\cdot\mathrm{erf}^{-1}(\alpha)$ (3)
@@
where $\mathrm{erf}^{-1}$ denotes inverse Gaussian error function. we can acquire the confidence interval $[\hat{y}_{\alpha/2},\hat{y}_{1-\alpha/2}]$, that satisfies $\mathcal{P}(y\in[\hat{y}_{\alpha/2},\hat{y}_{1-\alpha/2}])\geq 1-\alpha$.
@@
In practice, however, depression datasets [47, 48] exhibit high data imbalance, and the assumption of the Gaussian distribution may not be appropriate in this context. We prefer a distribution-free method for uncertainty quantification. Fortunately, with the introduction of pinball loss (Eq 4), we can train a DNN as a conditional quantile estimator [49, 50], allowing us to obtain arbitrary quantile values for $\mathcal{P}(y|X=x)$. Then, we can use ${\alpha/2}$ and ${1-\alpha/2}$ quantile values to derive the confidence interval $[\hat{y}_{\alpha/2},\hat{y}_{1-\alpha/2}]$, thereby achieving a distribution-free uncertainty quantification method.
Input: Calibration set $\mathcal{D}_{cal}=\{X_{i},y_{i}\}_{i=1}^{n}$, pre-trained depression recoginition model $\hat{f}(\cdot)$, test data $X_{n+1}$, miscoverage rate $\alpha\in(0,1)$, quantile estimator $\mathcal{Q}(\cdot,\cdot)$
Output: Confidence interval $\mathcal{C}_{\alpha}(X_{n+1})$
@@ 3      $\hat{y}_{i}=\hat{f}(X_{i})$
10 $\mathcal{C}_{\alpha}(X_{n+1})=[\hat{f}(X_{n+1})-\hat{s},\hat{f}(X_{n+1})+\hat{s}]$
return $\mathcal{C}_{\alpha}(X_{n+1})$
@@ 10 $\mathcal{C}_{\alpha}(X_{n+1})=[\hat{f}(X_{n+1})-\hat{s},\hat{f}(X_{n+1})+\hat{s}]$
The prerequisite for implementing the above method is that we can obtain a well-trained DNN model. However, due to limited depression datasets and imbalanced depression data distributions, using Eq 2 or Eq 4 to train depression recognition models is challenging. In practice, it has been observed that those methods suffer from overconfidence, where $\mathcal{P}(Y_{i}\in C(X_{i}))$ is significantly lower than $1-\alpha$. In this context, there is an urgent need for a post-processing method that is distribution-free to quantify the uncertainty of depression recognition. In the application of depression recognition, it’s impossible to access the conditional distribution $\mathcal{P}(y|X)$. To address this challenge, we can ensure that $\mathcal{C}(\cdot)$ guarantees marginal coverage to maintain an average coverage rate across the target, as expressed in Eq 5.
$\mathcal{P}(y_{test}\in\mathcal{C}(X_{test}))\geq 1-\alpha,\ \ \ (X_{test},y_{test})\in\mathcal{D}_{test}$ (5)
Conformal prediction (CP [27, 51]) addresses the requirement for uncertainty quantification in the field of depression recognition. We develop conformal depression prediction (CDP), which applies CP to the general depression recognition algorithm, as shown in Figure 2. The specific steps of CDP are shown in Algorithm 1. With CDP, we can acquire valid, theoretical coverage guaranteed confidence interval, and adjust the interval width according to user-defined confidence levels. We can ensure that this confidence interval achieves marginal coverage for the target depression scores at any confidence levels, as proven by [27, 51].
@@ III-C CDP with Approximate Conditional Coverage
There is a stronger notion 6 than the marginal coverage, namely conditional coverage.
$\mathcal{P}(y_{n+1}\in\mathcal{C}(X_{n+1})|X_{n+1})\geq 1-\alpha$ (6)
@@
for every input $X_{n+1}$, we aim to provide a confidence interval with $1-\alpha$ coverage. In contrast to marginal coverage, conditional coverage is more suitable for uncertainty quantification in individual depression prediction. In depression recognition applications, each individual only cares about whether the recognition prediction is accurate (or within which confidence interval does the depression score fall). Unfortunately, achieving conditional coverage is practically challenging in the finite datasets and a distribution-free setting [42]. Our work draws inspiration from ordinal adaptive prediction sets [26] in classification tasks. The motivation behind CDP-ACC is to generate adaptive confidence intervals that achieve approximate conditional coverage (Eq 6) in a model-agnostic and distribution-free manner for pretrained depression predictor.
Figure 3: The pipeline of CDP-ACC. Following the steps of CDP, CDP-ACC also requires a calibration set. We partition the predictions of the calibration set into bins. For samples within each bin, we estimate the distribution of depression scores $y$ using histograms, and calculate the conformal scores within the group according to Eq 7.
@@
Approximate Conditional Coverage. In pretrained depression predictor, $\hat{y}=\mathbb{E}[\mathcal{P}(y|X)]$, we cannot directly access to the distribution $\mathcal{P}(y|X)$. To accurately estimate $\mathcal{P}(y|X)$ in depression recognition, we rely on a key assumption from [52]: similar $X$ will have similar conditional distributions $\mathcal{P}(y|X)$. Formally, if $X_{i}\approx X_{j}$, $\mathcal{P}(y|X=X_{i})\approx\mathcal{P}(y|X=X_{j})$, that is $\mathbb{E}[\mathcal{P}(y|X_{i})]\approx\mathbb{E}[\mathcal{P}(y|X_{j})]$. This assumption is also reasonable in our practical observations. For example, different video clips of the same patient tend to have similar depression predictions.
@@
Then, suppose there are samples $X$ belonging to a very small neighborhood $\Delta(X_{i})$, which we can consider they share the similar conditional distribution according to the smooth assumption above. The straightforward idea to approximate $\mathcal{P}(y|X=X_{i})$ is to replace $X=X_{i}$ with $X\in\Delta(X_{i})$, that is $\mathcal{P}(y|X=X_{i})\approx\mathcal{P}(y|X\in\Delta(X_{i}))$. In the context of facial depression recognition, $X_{i}$ denotes facial video clips. $\Delta(X_{i})$ can represent clips belonging to the same video, or clips with the same depression score. However, for unseen test video clips, we have no oracle with access to $\Delta(X_{i})$. We need to further explore relaxation to approximate $\Delta(X_{i})$.
@@
We revisit $\Delta(X_{i})$ from the perspective of the model (pre-trained model $\hat{f}$): $\Delta(\hat{y_{i}})=\hat{f}(\Delta(X_{i}))$. The model predicts similar depression scores for what it perceives as similar $X_{i}$. Thus, we can use $\Delta(\hat{y_{i}})$ as the condition instead of $\Delta(X_{i})$. That is, we further assume based on [52]: similar $\hat{y}$ also have similar conditional distributions [^1] This assumption can be understood as follows: we all know that the mean squared error (MSE) loss can be derived as the maximum likelihood estimate of a Gaussian distribution with a mean of $\hat{y}$ and a fixed variance. Therefore, if $\hat{y}$ is similar, then the fixed-variance Gaussian distribution will be similar.. Then, $\mathcal{P}(y|X=X_{i})\approx\mathcal{P}(y|\hat{y}\in\Delta(\hat{y}))$. Given the smoothness assumption (demo shown in Figure 4), we can estimate the conditional distribution by grouping neighboring points of the depression predictions.
Figure 4: Demo of smoothed assumption. We measure the uncertainty of the model predictions by considering $\mathcal{P}(y|\hat{y}\in\Delta(\hat{y})$. For example, we consider all individuals with depression predictions $\hat{y}=1\sim 3$ as a group. If the model’s predictions are reliable, the actual depression scores for this group would be about $1\sim 3$. If the model’s predictions are unreliable, then there will be abnormal depression scores in this group.
@@
Conformal score. In conformal prediction, the setting of conformal score is particularly important. Different conformal scores essentially determine the effectiveness of the conformal method. In CDP-ACC, the setting of conformal score consists of dividing bins and constructing a nested sequence of approximate oracle intervals.
In $\mathcal{D}_{cal}$, we first divide $\hat{y}\in[0,K)$ into $M$ bins: $B_{m}=[\frac{m-1}{M}K,\frac{m}{M}K)$, where $m=1,...,M$ and $K$ denotes the predetermined upper bound. Now, we have $M$ bins and predictions $\hat{y}\in B_{m}$ fall in bins. Based on the approximation $\mathcal{P}(y|X=X_{i})\approx\mathcal{P}(y|\hat{y}\in B_{m})$, we only require the depression label $y$ corresponding to $\hat{y}$ within $B_{m}$ to approximately obtain the conditional distribution $\mathcal{P}(y|X=X_{i})$.
Subsequently, we need to estimate the distribution of $y$ within each bin. Referring to [51, 26], we use histogram estimation to acquire $\mathcal{P}_{m}(y|\hat{y}\in B_{m})$ for bin $m$. A good confidence interval should achieve the specified coverage rate. Narrower interval widths are desirable as they provide more precise uncertainty quantification.
@@
To maintain the continuity of prediction interval, we construct a nested sequence $E(\cdot)$ so that the interval with higher confidence covers the interval with lower confidence, all while endeavoring to minimize the width of each interval, as shown in Figure 5.
$E_{m}(\tau,\mathcal{P}_{m})=\operatorname*{argmin}_{[y_{lo},y_{hi}]\subseteq{B_{m}}}\{y_{hi}-y_{lo}:\mathcal{P}_{m}(y\in[y_{lo},y_{hi}])\geq\tau\}$ (7)
@@
where $\tau\in(0,1)$ and $m\in[1,M]$. If the optimal solution is not unique, we can add a little noise into the $B_{m}$ to break the ties. It should note that $E_{m}(\cdot)$ is a nested sequence for bin $m$, for example $E_{m}(0.1)\subseteq E_{m}(0.2)\subseteq E_{m}(0.3)$, and $E_{m}(\tau)$ is a interval $[y_{m,lo},y_{m,hi}]$.
Figure 5: The specific process of conformal score. First, bins are divided according to $\hat{y}$, and we use each bin to denote the neighborhood of $\hat{y}$. Second, we compute the sub-conformal-score by histogram estimation within each bin. Finally, the predicted interval can be acquired based on the bin assigned to the test sample $\hat{f}(X_{n+1})$.
@@
Confidence Interval. After calculating the conformal scores in the calibration set, we can leverage the assumption that the calibration set and the test samples are exchangeable to generate valid confidence intervals. Specifically, given all conformal scores $E_{m}$ in the calibration set, a user-defined miscoverage rate $\alpha$, and an unseen test sample $X_{n+1}$, CDP-ACC acquires confidence intervals according to the following steps. First, for each bin $m$, calculate the $(1-\alpha)$th quantile of the conformal scores using Eq. 7, and acquire the corresponding interval $[y_{m,lo},y_{m,hi}]$. Within each bin of the calibration set, there exists a confidence interval with a coverage rate of $1-\alpha$ and the shortest width. Then, use the pre-trained model $\hat{f}$ to predict the depression score of $X_{n+1}$, i.e., $\hat{y}=\hat{f}(X_{n+1})$. The corresponding confidence interval $[y_{m,lo},y_{m,hi}]$ can be obtained based on the bin $m$ into which $\hat{y}_{n+1}$ falls.
Input: Calibration set $\mathcal{D}_{cal}=\{X_{i},y_{i}\}_{i=1}^{n}$, pre-trained depression recoginition model $\hat{f}(\cdot)$, test data $X_{n+1}$, miscoverage rate $\alpha\in(0,1)$, upper bound $K$, histogram estimator $\mathcal{H}(\cdot)$, bins number $M$
@@
Output: Confidence interval $\mathcal{C}_{\alpha}(X_{n+1})$
2 Initialize bins $\{B_{m}\}_{m=0}^{M-1}\leftarrow\emptyset$
// Bins dividing for $\hat{y}$ on $\mathcal{D}_{cal}$

### v3

@@ III-B Depression Prediction with UQ
The straightforward method for uncertainty quantification of depression prediction is to train a DNN to estimate the conditional probability distribution $P(y|X=x)$ of depressed individuals. Given our lack of knowledge regarding the true $P(y|X=x)$, one possible solution is to assume that the depression target $y$ follow a Gaussian distribution $y\sim\mathcal{N}(\hat{y},\sigma^{2})$, and uses the negative log-likelihood (NLL) [47] as the loss function:
$\mathcal{L}_{NLL}=\sum_{i=1}^{|\mathcal{D}_{train}|}\frac{\text{log}\sigma_{i}^{2}}{2}+\frac{(y_{i}-\hat{y}_{i})^{2}}{2\sigma_{i}^{2}}$ (2)
@@
The learning objective allows us to train a DNN capable of predicting the distribution $P(y|X=x)$ for any given individual $x$. Furthermore, under the Gaussian assumption, the $\alpha$-th quantile can be calculated by
$\hat{y}_{\alpha}=\hat{y}+\sqrt{2}\sigma\cdot\mathrm{erf}^{-1}(2\alpha-1)$ (3)
@@
where $\mathrm{erf}^{-1}$ denotes the inverse Gaussian error function. In this way, we can obtain the confidence interval $[\hat{y}_{\alpha/2},\hat{y}_{1-\alpha/2}]$ that satisfies $P(y\in[\hat{y}_{\alpha/2},\hat{y}_{1-\alpha/2}])\geq 1-\alpha$.
@@
In practice, however, depression datasets (e.g., AVEC 2013 and AVEC 2014 [48, 49]) exhibit high data imbalance, and the assumption of the Gaussian distribution may not be appropriate in this context. We prefer a distribution-free method for uncertainty quantification. Indeed, by introducing the pinball loss (Eq. (4)), we can train a DNN as a conditional quantile estimator [50, 51], allowing us to obtain arbitrary quantile values for $P(y|X=x)$:
where $q\in[0,1]$, $\hat{y}_{q}$ denotes the $q$-th quantile of the prediction, and $y$ is the ground truth. Using the trained regressor, we can use ${\alpha/2}$ and ${1-\alpha/2}$ quantile values of the predictive distribution to derive the confidence interval $[\hat{y}_{\alpha/2},\hat{y}_{1-\alpha/2}]$, thereby achieving a distribution-free uncertainty quantification.
Input: $\mathcal{D}_{cal}=\{X_{i},y_{i}\}_{i=1}^{N}$: calibration set;
$\hat{f}(\cdot)$: pre-trained depression prediction model;
$X_{test}$: test data;
$\alpha\in(0,1)$: miscoverage rate;
$\mathcal{Q}(\cdot,\cdot)$: quantile estimator
Output: Confidence interval $\mathcal{C}_{\alpha}(X_{test})$
@@ 3      $\hat{y}_{i}=\hat{f}(X_{i})$
10 $\mathcal{C}_{\alpha}(X_{test})=[\hat{f}(X_{test})-\hat{s},\hat{f}(X_{test})+\hat{s}]$
return $\mathcal{C}_{\alpha}(X_{test})$
@@ 10 $\mathcal{C}_{\alpha}(X_{test})=[\hat{f}(X_{test})-\hat{s},\hat{f}(X_{test})+\hat{s}]$
However, due to limited depression data and imbalanced data distributions, using Eq. (2) or Eq. (4) to train depression prediction models is challenging. In practice, it has been observed that those methods suffer from overconfidence, where $P(y\in\mathcal{C}_{\alpha}(X))$ can be significantly lower than $1-\alpha$. In this context, there is an urgent need for a post-processing method that is distribution-free to quantify the uncertainty of depression prediction. In particular, we seek for constructing the confidence interval $\mathcal{C}_{\alpha}(\cdot)$ to guarantee the marginal coverage so as to maintain an average coverage rate across the target, as expressed by
$P(y_{test}\in\mathcal{C}_{\alpha}(X_{test}))\geq 1-\alpha,\ \ \ (X_{test},y_{test})\in\mathcal{D}_{test}$ (5)
To that end, we introduce conformal prediction (CP [26, 52]) for uncertainty quantification of facial depression prediction. Specifically, we develop conformal depression prediction (CDP) as shown in Fig. 2. As premise of CP is to have a calibration set that is exchangeable with the test sample (This is guaranteed if the i.i.d. assumption holds), we calculate the conformal score $s$ on the calibration set,
$s_{i}=|y_{i}-\hat{y}_{i}|$ (6)
which reflects the discrepancy between the depression prediction and the ground truth. Next, we calculate the $q$-th quantile of $s$ and denote it as $\hat{s}$. For any test sample $X_{test}$, we have $P(s_{test}\leq\hat{s})=q$. In this way, we can obtain its prediction intervals $\mathcal{C}_{\alpha}(X_{test})=[\hat{y}_{test}-\hat{s},\hat{y}_{test}+\hat{s}]$, which is valid with
$1-\alpha\leq P(y_{test}\in\mathcal{C}_{\alpha}(X_{test})\leq 1-\alpha+\frac{1}{N+1}$ (7)
The implementation of CDP are summarized in Algorithm 1. With CDP, we can acquire valid, theoretical coverage guaranteed confidence interval, and the interval width relates to the user-defined confidence levels $1-\alpha$. Note that this CP-based uncertainty quantification achieves marginal coverage for the target depression scores at any confidence levels, as proven in [26, 52].
@@ III-C CDP with Approximate Conditional Coverage
There is a stronger notion than the marginal coverage, namely conditional coverage, as defined by
$P(y_{test}\in\mathcal{C}_{\alpha}(y|X_{test}))\geq 1-\alpha$ (8)
@@
For the model prediction $\hat{y}_{test}$ on any test sample $X_{test}$, it aims to provide a confidence interval $C_{\alpha}$ with $1-\alpha$ coverage. In contrast to marginal coverage, conditional coverage is more suitable for uncertainty quantification in individual depression prediction. In practical applications, each individual only cares about whether the prediction is trustworthy (or within which confidence interval the prediction falls). Unfortunately, achieving conditional coverage is practically challenging in a distribution-free setting [44]. Our work draws inspiration from ordinal adaptive prediction sets [53] for classification tasks. The motivation behind our CDP-ACC is to generate adaptive confidence intervals for pre-trained depression predictor, achieving approximate conditional coverage in a model-agnostic and distribution-free manner.
Figure 3: The pipeline of CDP-ACC. The predictions on the calibration set are firstly partitioned into $M$ subintervals. For each subinterval, CDP-ACC estimates the approximate conditional distribution $P(y|\hat{y})$, computes the conformal score based on the calibration samples in that interval, and finally constructs the adaptive confidence interval.
III-C1 Approximate Conditional Coverage
@@
Given a pre-trained depression predictor with point prediction, $\hat{y}=\hat{f}(X)$, we cannot directly access to the conditional distribution $P(y|X)$. To accurately estimate $P(y|X)$ in depression prediction, we rely on a key assumption [54]: similar $X$ (in feature embedding space) will have similar conditional distributions $P(y|X)$ and similar conditional expectation. This assumption is reasonable in our practical observations. For example, different video clips of the same patient tend to have similar depression predictions.
@@
Let $\Delta(X_{i})$ denote the neighborhood of $X_{i}$ in the feature embedding space. Our aim is to approximate $P(y|X=X_{i})$ with $P(y|X\in\Delta(X_{i}))$. In the context of facial depression prediction, $X_{i}$ denotes the facial video clips, and $\Delta(X_{i})$ can represent clips belonging to the same video, or clips with the same depression score. However, for unseen test video clips, we do not have oracle with access to $\Delta(X_{i})$. We need to further explore relaxation scheme to approximate $\Delta(X_{i})$.
@@
We revisit $\Delta(X_{i})$ from the perspective of the model predictions (with pre-trained model $\hat{f}$): $\Delta(\hat{y_{i}})=\hat{f}(\Delta(X_{i}))$. The model predicts similar depression scores for what it perceives as similar $X_{i}$. Thus, we can use $\Delta(\hat{y_{i}})$ as the condition instead of $\Delta(X_{i})$, that is, similar predictions $\hat{y}$ can also have similar conditional distributions. With this approximation, we have $P(y|X=X_{i})\approx P(y|\hat{y}\in\Delta(\hat{y}))$, and we can estimate the conditional distribution by grouping neighboring points of the depression predictions. Moreover, by incorporating the approximation, we can introduce inductive bias, which implies that for samples falling in $\Delta(\hat{y})$, the more concentrated the points, the higher the credibility.
We uniformly divide the range of $\hat{y}$ into $M$ subintervals $B_{m},\ m=1,...,M$. Let $y_{m,L}$ and $y_{m,U}$ denote the lower and upper bounds of $y$ in each subinterval $m$. Based on the approximation $P(y|X=X_{i})\approx P(y|\hat{y}\in B_{m})$, we can only use the depression label $y$ (corresponding to $\hat{y}$ within $B_{m}$) to approximately estimate the conditional distribution $P(y|X=X_{i})$. Following [52, 53], we use histogram binning to estimate the conditional probability distribution $P_{m}(y|\hat{y}\in B_{m})$ for each subinterval $m$ (shown in Fig. 4).
Figure 4: Histogram estimation of the conditional distribution $P(y|\hat{y}\in\Delta(\hat{y}))$. For samples falling in $\hat{y}\in\Delta(\hat{y})$, the more concentrated the points, the higher the credibility.
III-C2 Confidence Interval
@@
In conformal prediction, setting the conformal score is particularly important. The choice conformal score function essentially determines the effectiveness of the conformal prediction method. In CDP-ACC, setting the conformal score involves constructing nested sequence of approximate oracle intervals.
@@
To maintain the continuity of prediction interval, we construct a nested sequence $E(\cdot)$ so that the interval with higher confidence covers the interval with lower confidence, all while endeavoring to minimize the width of each interval. As shown in Fig. 5, for each $y_{i}$ whose prediction falls into $B_{m}$ we can construct the confidence interval $E_{m}(\tau,y_{i})$ with the shortest width:
$E_{m}(\tau,y_{i})=\mathop{\arg\min}\limits_{\begin{subarray}{c}[y_{l},y_{h}]\subset[y_{m,L},y_{m,U}],\\ y_{i}\in[y_{l},y_{h}]\end{subarray}}$$\{y_{h}-y_{l}:$ (9)
$F_{m}(y_{h}|\hat{y})-F_{m}(y_{l}|\hat{y})\geq\tau\}$
@@
where $\tau\in(0,1)$ denotes the confidence level, $m\in[1,M]$ and $F_{m}$ denotes the cumulative probability distribution of the conditional probability distribution $P_{m}(y|\hat{y}\in B_{m})$. If the optimal solution is not unique, we can add a little noise $\epsilon$ into the $B_{m}$ to break the ties [55]. It should be noted that, given $y_{i}$, $E_{m}(\tau,y_{i})$ is a nested sequence as $\tau$ grows, for example, $E_{m}(0.1,y_{i})\subseteq E_{m}(0.2,y_{i})\subseteq E_{m}(0.3,y_{i})$. For any sample $y_{i}$ whose prediction falls into $B_{m}$, we set the width of its interval $E_{m}$ as the conformal score $s_{i}$:
$s_{i}(m,\alpha)=|E_{m}(1-\alpha,y_{i})|$ (10)
The conformal score typically reflects the prediction bias. Higher scores indicate poorer prediction quality while wider interval corresponds to worse predictions in our context.
Figure 5: The setting of the conformal score for CDP-ACC. Given the approximate conditional distribution $P(y|\hat{y})$, various prediction intervals can be acquired containing $y_{i}$ with $1-\alpha$ coverage, of which the width of the shortest one is set as the conformal score for our CDP-ACC.
@@
After calculating the conformal scores on the calibration set, we can estimate the confidence interval for any test samples based on the assumption that the calibration set and the test samples are exchangeable to generate valid confidence intervals. A good confidence interval should achieve the specified coverage rate at least. Narrower interval widths are desirable as they provide more precise uncertainty quantification. For $E_{m}(\tau,y_{i})$, if the user sets the confidence level $\tau$ to 100%, the confidence interval for any samples $y_{i}$ will be unique ($E_{m}(1,y_{i})$). Otherwise, there will be multiple prediction intervals that cover $y_{i}$ with at least $1-\alpha$ probability. According to our smoothness approximation, the concentration degree of samples within a subinterval can represent prediction uncertainty. We use interval width as the conformal score and identify the narrowest interval that covers $y_{i}$ with $1-\alpha$ probability as its confidence prediction interval. Specifically, given all conformal scores calculated on the calibration set and a user-defined miscoverage rate $\alpha$, CDP-ACC constructs confidence interval for any unseen test sample $X_{test}$ according to the following steps. First, for each subinterval $m$, we calculate the $(1-\alpha)$-th quantile of the conformal scores $\{s_{i}(m,\alpha)\}$.
$\hat{s}(m,\alpha)=\mathcal{Q}(\{s_{i}(m,\alpha)\},1-\alpha)$ (11)
Then, for each subinterval $m$ we can obtain its confidence interval with shortest width according to
$\mathcal{A}(m,\alpha)=\mathop{\arg\min}\limits_{\begin{subarray}{c}[y_{l},y_{h}]\subset[y_{m,L},y_{m,U}],\\ y_{i}\in[y_{l},y_{h}]\end{subarray}}\{y_{h}-y_{l}:s_{i}(m,\alpha)\leq\hat{s}(m,\alpha)\}$ (12)
Finally, based on the interval into which the prediction of the test sample falls, we can retrieve the corresponding confidence interval. The proposed CDP-ACC is summarized in Algorithm 2.
Input: $\mathcal{D}_{cal}=\{X_{i},y_{i}\}_{i=1}^{N}$: calibration set;
$\hat{f}(\cdot)$: pre-trained depression prediction model;
$X_{test}$: test data;
$\alpha\in(0,1)$: miscoverage rate;
$L,U$: lower and upper bounds of the predictions on $\mathcal{D}_{cal}$;
$\mathcal{H}(\cdot)$: histogram estimator;
$\mathcal{Q}(\cdot,\cdot)$: quantile estimator;
$M$: number of the subintervals
@@
Output: Confidence interval $\mathcal{C}_{\alpha}(X_{test})$
2 Initialize subintervals $\{B_{m}\}_{m=1}^{M}\leftarrow\emptyset$
// Subintervals dividing for $\hat{y}$ on $\mathcal{D}_{cal}$